○student population of a university consists of 30% asians, 40% americans, 20% europeans, 10% africans. It is known that 40% of all asian students, 50% of all american students, 60% of all european students and 20% of all african students are girls. Find the probability that a girl chosen at random from the university is an Asian.
1
Expert's answer
2020-06-22T18:14:44-0400

Following contingency table is formed from given data:


Gender/Race Asians Americans European African Total

Girl 0.4*30= 12 0.5*40=20 0.6*20=12 0.2*10=2 46

Boy 18 20 8 8 54

total 30 40 20 10 100


P(girl who is Asian)=12/100=0.12